1. Express the product Ax as a linear combination of the column vectors of A.
Close and open parenthesis 4 1 Close and open Parenthesis 2
-2 3 3

2. Suppose that the solution set of the homogeneous system Ax = 0 is given by the formulas

x1 = − 5r + 6s, x2 = r − 3s, x3 = r, x4 = s

Find a vector form of the general solution of Ax = 0.

3. Suppose that x1 = − 3, x2 = 4, x3 = 1, x4 = − 2 is a solution of a non homogeneous linear system Ax = b, and that the solution set of the homogeneous system Ax = 0 is given by the formulas x1 = − 5r + 6s, x2 = r − 4s, x3 = r, x4 = s. Find a vector form of the general solution of Ax = b.
1
Expert's answer
2020-01-07T09:50:01-0500

1.


"A=\begin{pmatrix}
 2 & -2 \\
 3 & 3
\end{pmatrix}, x=\begin{pmatrix}
 4 \\
 1 
\end{pmatrix}"

"Ax=\begin{pmatrix}
 2 & -2 \\
 3 & 3
\end{pmatrix}\begin{pmatrix}
 4 \\
 1 
\end{pmatrix}=\begin{pmatrix}
 2\cdot4+(-2)\cdot1 \\
 3\cdot4+3\cdot1 
\end{pmatrix}="

"=4\begin{pmatrix}
 2 \\
 3
\end{pmatrix}+1\begin{pmatrix}
 -2 \\
 3 
\end{pmatrix}"

2.

The solution set of the homogeneous system Ax = 0 is given by the formulas


"x_1=-5r+6s, x_2=r-3s,x_3=r,x_4=s"

The solution can be written in matrix form. Then


"\begin{bmatrix}
 x_1 \\
 x_2 \\
 x_3 \\
 x_4 
\end{bmatrix}=\begin{bmatrix}
 -5r+6s \\
 r-3s \\
 r \\
 s 
\end{bmatrix}=r\begin{bmatrix}
 -5 \\
 1 \\
 1 \\
 0 
\end{bmatrix}+s\begin{bmatrix}
 6 \\
 -3 \\
 0 \\
 1 
\end{bmatrix}"

Then


"x=r\begin{bmatrix}
 -5 \\
 1 \\
 1 \\
 0 
\end{bmatrix}+s\begin{bmatrix}
 6 \\
 -3 \\
 0 \\
 1 
\end{bmatrix}"

3.

The solution set of the homogeneous system Ax = 0 is given by the formulas


"x_1=-5r+6s, x_2=r-4s,x_3=r,x_4=s"

The solution can be written in matrix form. Then


"\begin{bmatrix}
 x_1 \\
 x_2 \\
 x_3 \\
 x_4 
\end{bmatrix}=\begin{bmatrix}
 -5r+6s \\
 r-4s \\
 r \\
 s 
\end{bmatrix}=r\begin{bmatrix}
 -5 \\
 1 \\
 1 \\
 0 
\end{bmatrix}+s\begin{bmatrix}
 6 \\
 -4 \\
 0 \\
 1 
\end{bmatrix}"

The solution set of the non homogeneous system Ax = b is given by the formulas


"x_1=-3, x_2=4,x_3=1,x_4=-2"

The solution can be written in matrix form. Then


"\begin{bmatrix}
 x_1 \\
 x_2 \\
 x_3 \\
 x_4 
\end{bmatrix}=\begin{bmatrix}
 -3 \\
 4 \\
 1 \\
 -2
\end{bmatrix}"



Hence the general solution is


"x=\begin{bmatrix}
 -3 \\
 4 \\
 1 \\
 -2
\end{bmatrix}+r\begin{bmatrix}
 -5 \\
 1 \\
 1 \\
 0 
\end{bmatrix}+s\begin{bmatrix}
 6 \\
 -4 \\
 0 \\
 1 
\end{bmatrix}"